"\int_{y=1}^3\int_{x=0}^{3}" (1 + 8xy) dx dy


1
Expert's answer
2021-07-26T11:11:40-0400
"\displaystyle\int_{y=1}^3\displaystyle\int_{x=0}^3(1+8xy)dxdy"

"=\displaystyle\int_{y=1}^3\big[x+4x^2y\big]\begin{matrix}
 3 \\
 0
\end{matrix}dy"

"=\displaystyle\int_{y=1}^3(3+36y)dy"

"=\big[3y+18y^2\big]\begin{matrix}
 3 \\
 1
\end{matrix}"

"=9+162-3-18=150"